OpenStudy (anonymous):

The hypotenuse of a right triangle is 50 mm long. One leg of the right triangle is 30 mm long. What is the length of the other leg?

jimthompson5910 (jim_thompson5910):

a^2+b^2=c^2 30^2+b^2 = 50^2 900+b^2=2500 b^2 = 2500-900 b^2 = 1600 b = sqrt(1600) b = 40 So the length of the other leg is 40 mm

jimthompson5910 (jim_thompson5910):

To get to the answer quicker, notice that this 30, 40, 50 triangle has lengths that correspond to a 3, 4, 5 triangle

OpenStudy (anonymous):

your officially my savior :P